# Fast Tree-Field Integrators: From Low Displacement Rank to Topological Transformers

### v1

@@ 4.4 Large Scale Transformer Experiments using FTFI
We run experiments on ImageNet and Places365 datasets using ViT-B/16 (see Table 1). For all the kernels, our variants beat the baselines. For $\phi(x)=x^{4}$, the best variant applies an exponentiated quadratic function, for which we apply Vandermonde matrices (see: discussion in Sec. 3.2.1). Our best variant across all kernels (78.79%) provides 2% accuracy gains over the best baseline (76.76%). In the synced setting, we use only three extra learnable parameters per layer (shared in all attention heads across all layers) and obtain 1-1.5% accuracy gains. In the asynced setting, we use a small set of $\mathbf{36}$ extra learnable parameters per layer (3 extra parameters per head). Overall, we observe that FTFI improves the approximation quality within Transformers with a minimal number of parameters.
@@
We provide additional experiments including Gromov-Wasserstein distance computation (Vayer et al., 2018) (see Sec. D.2), along with code pointers (Appendix D).
@@ Appendix A Theoretical results
We will apply Lemma 7.19 from (Cygan et al., 2015) (that we provide also below for reader’s convenience) and its algorithmic proof. We refer to Cygan et al. (2015) for a definition of the related graph terms.
@@ A.2.2 Non-Uniform FFT (NU-FFT)
We will now propose a closely-related method, that relies on the non-uniform FFT (NU-FFT). [^1] See (Greengard and Lee, 2004) for an excellent introduction.
@@
So far we have not put any restrictions on the tree weights. If we restrict all weights to be the same (without loss of generality, equal to one), then the problem becomes easier. In this case for any function $f$, matrices $\mathbf{C}$ and $\mathbf{C}^{\top}$ are Hankel (Brent, 2010) (constant on each anti-diagonal and belonging to LDR class). Then, matrix-vector multiplication can be done in $O((a+b)\log(a+b))$. The analysis from the proof of Lemma 3.3 for $d=1$ can be repeated. We conclude that $f$-integration can be conducted in $O(N\log^{2}(N))$ time for $N$-vertex unweighted trees and any $f:\mathbb{R}\rightarrow\mathbb{R}$. This was already proven in (Choromanski et al., 2022).
@@ Appendix B Additional Related Work
In this section we provide additional related works. One of the methods to tackle this problem is via iterative methods (Koutis et al., 2012) like Arnoldi iteration (Arnoldi, 1951), Conjugate Gradient (Shewchuk, 1994) and the celebrated Spielman-Teng algorithm (Spielman and Teng, 2012) for symmetric diagonally dominant (SDD) matrices. There are a number of extensions and variations of the above methods (Blelloch et al., 2011; Boman et al., 2008; Christiano et al., 2010; Koutis and Miller, 2007; Spielman and Teng, 2008; Daitch and Spielman, 2008; Koutis and Miller, 2008).They mainly take into account the structure of the matrix (SDD) (Koutis et al., 2010, 2011a, 2012), embedding of a graph into low stretch spanning trees (Elkin et al., 2005), graph sparsification (Spielman and Teng, 2010) and the choice of a good pre-conditioner (Maggs et al., 2003; Koutis et al., 2011b). We want to emphasize that the research on low stretch trees for general graphs is orthogonal to the main topic of this work. In our manuscript, we show in particular how to conduct efficient integration on arbitrary trees. Thus our work can be naturally combined with those algorithms to leverage all the above low stretch tree constructions for a better approximation of the graph’s metric.
The other class of method comes from the celebrated work of (Al-Mohy and Higham, 2011) and there are a number of extensions of this work (Kloster and Gleich, 2023; Al-Mohy and Higham, 2010; Moore, 2011; Moler and Van Loan, 2003; Auckenthaler et al., 2010).
Another class of methods is via sampling, where one samples a subset of a large matrix, which is then used to approximate the matrix-vector multiplication (i.e. Monte Carlo sampling) methods (Drineas et al., 2006; Drineas and Kannan, 2001; Acebron, 2019; Acebron et al., 2019; Benzi et al., 2017; Martinsson, 2019).
@@ Appendix C Topological Transformers
Algorithm 1 General Efficient Low-Rank Masked Attention from Choromanski et al. (2022)
@@
The central question in (Choromanski et al., 2022) was how to incorporate the masking in the linear attention as above. Note that in this case $\mathbf{A^{\prime}}$ is never materialized. Building on the work of (Luo et al., 2021), the authors (Choromanski et al., 2022) propose a general algorithm that efficiently implements masked linear attention.
@@ Appendix D Experimental Details and Additional Experiments
In this section, we provide additional details regarding the experimental setup and present additional results from our experiments. We will open source all our code at the time of publication. Specifically, we provide there the code for: (1) our algorithm leveraging IntegratorTree data structure (depicted in Fig 1), (2) adaptation to the Gromov-Wasserstein-type computation, (3) graph classification and (4) experiments on interpolation on meshes.
@@ D.1 Additional experiments for graph metric approximation with $f$-distance matrices
We also perform similar experiments for graph classification on the CUBES dataset Hanocka et al. (2019). Specifically, we investigate how the polynomial degree affects the graph classification performance in Fig. 9 (left). We observe that increasing the polynomial degree improves the classification accuracy up to a certain degree. For the same dataset, we also compute the training loss for different polynomial degrees in Fig. 9 (right). Similarly, we observe that higher-degree rational functions achieve lower training loss for fitting the polynomial coefficients.
@@
Moreover, we benchmark FTFI on ModelNet10 (Wu et al., 2015), a dataset for 3D Point Cloud (PC) classification. For each PC, we create an $\epsilon$-neighborhood-graph and use FTFI for graph classification The Shortest Path kernel achieves an accuracy of $39.6\%$, whereas our FTFI with the degree-2 polynomial improves the accuracy to $44.2$% ($10$% relative improvement over the baseline), similar to the observation in 9.
@@ D.2 Integration of FTFI into GW-style algorithms
Wasserstein distance has found many uses in ML, particularly due to it’s principled approach to compare probability distributions. Gromov Wasserstein Mémoli (2011) discrepancy is an extension of Wasserstein distance to graph structured data, with a lot of downstream applications like graph clustering and classification. Inspired by the work of (Choromanski et al., 2023), we follow the exact same procedure in the integration of FTFI in the conditional gradient algorithm. The FTFI can be injected seamlessly in place of the Fast Matrix Multiplication (FMM) algorithms in Algorithm 2 and Algorithm 3 (see (Choromanski et al., 2023)).
Our method GW-FTFI run consistently $2$-$6$x faster than the baseline methods using the Shortest Path kernel, with no drop in accuracy in computing the associated costs (Figure 10). The plots shown are obtained by averaging over $10$ seeds and random trees of various sizes. For the baseline experiments, we use the implementation from the POT library (Flamary et al., 2021).
@@ D.3 Interpolation on Meshes
In the vertex normal prediction task in Section 4.2, we choose 40 meshes for 3D-printed objects with a wide range of size from the Thingi10K (Zhou and Jacobson, 2016) dataset with the File IDs:
@@ D.4 Additional Details on Graph Classification
We conduct graph classification experiments on a wide range of benchmark datasets. We report the dataset statistics for the graph classification datasets in Table 2. More details about the datasets are available in Morris et al. (2020). To evaluate the performance of the different kernels, we employ the framework proposed by (Errica et al., 2020). In particular, 10-fold cross-validation is used to obtain an estimate of the generalization performance of our method and the baseline method. We repeat this cross validation experiment 5 times to get a robust estimation and report the standard deviation for each setup.
To obtain graph features, we follow the approach presented in (de Lara and Pineau, 2018). In this setting, we obtain the $k$-smallest eigenvalues from the approximated kernel from FTFI and forward these features to a random forest classifier for classification. For BGFI, we perform the same process obtaining the $k$-smallest eigenvalues from the exact shortest kernel. FTFI achieves similar performance to the BGFI while being significantly faster. We tune the hyperparameter $k$ independently for each method.
@@
Table 4: Comparison of FTFI with a broad range of graph kernel-based classification approaches. We observe that FTFI achieves performance similar to that of Exact SP, its exact counterpart, across almost all datasets. The baseline results have been compiled from Nikolentzos et al. (2021). OOT and OOM indicate that the corresponding algorithm ran out of time or memory respectively.

### v2

@@ 4.4 Large Scale Transformer Experiments using FTFI
We run experiments on ImageNet and Places365 datasets using ViT-B/16 (see Table 1). For all the kernels, our variants beat the baselines. For $\phi(x)=x^{4}$, the best variant applies an exponentiated quadratic function, for which we apply Vandermonde matrices (see: discussion in Sec. 3.2.1). Our best variant across all kernels (78.79%) provides 2% accuracy gains over the best baseline (76.76%). In the synced setting, we use only three extra learnable parameters per layer (shared in all attention heads across all layers) and obtain 1-1.5% accuracy gains. In the asynced setting, we use a small set of $\mathbf{36}$ extra learnable parameters per layer (3 extra parameters per head). Overall, we observe that FTFI improves the approximation quality within Transformers with a minimal number of parameters. We provide additional discussions on the ViT results for ImageNet in Appendix D.5.1 and for Places365 in Appendix D.5.2.
Additional results on the I-Naturalist dataset, where we outperform various low-rank attention baselines, are provided in Appendix D.5.3.
@@
Further results on Video Transformer (ViViT) (Arnab et al., 2021) are provided in Appendix D.6. We also provide additional experiments including Gromov-Wasserstein distance computation (Vayer et al., 2018) (see Sec. D.2), along with code pointers (Appendix D).
@@ Appendix A Theoretical results
We will apply Lemma 7.19 from [Cygan et al., 2015] (that we provide also below for reader’s convenience) and its algorithmic proof. We refer to Cygan et al. [2015] for a definition of the related graph terms.
@@ A.2.2 Non-Uniform FFT (NU-FFT)
We will now propose a closely-related method, that relies on the non-uniform FFT (NU-FFT). [^1] See [Greengard and Lee, 2004] for an excellent introduction.
@@
So far we have not put any restrictions on the tree weights. If we restrict all weights to be the same (without loss of generality, equal to one), then the problem becomes easier. In this case for any function $f$, matrices $\mathbf{C}$ and $\mathbf{C}^{\top}$ are Hankel [Brent, 2010] (constant on each anti-diagonal and belonging to LDR class). Then, matrix-vector multiplication can be done in $O((a+b)\log(a+b))$. The analysis from the proof of Lemma 3.3 for $d=1$ can be repeated. We conclude that $f$-integration can be conducted in $O(N\log^{2}(N))$ time for $N$-vertex unweighted trees and any $f:\mathbb{R}\rightarrow\mathbb{R}$. This was already proven in [Choromanski et al., 2022].
@@ Appendix B Additional Related Work
In this section we provide additional related works. One of the methods to tackle this problem is via iterative methods [Koutis et al., 2012] like Arnoldi iteration [Arnoldi, 1951], Conjugate Gradient [Shewchuk, 1994] and the celebrated Spielman-Teng algorithm [Spielman and Teng, 2012] for symmetric diagonally dominant (SDD) matrices. There are a number of extensions and variations of the above methods [Blelloch et al., 2011, Boman et al., 2008, Christiano et al., 2010, Koutis and Miller, 2007, Spielman and Teng, 2008, Daitch and Spielman, 2008, Koutis and Miller, 2008].They mainly take into account the structure of the matrix (SDD) [Koutis et al., 2010, 2011a, 2012], embedding of a graph into low stretch spanning trees [Elkin et al., 2005], graph sparsification [Spielman and Teng, 2010] and the choice of a good pre-conditioner [Maggs et al., 2003, Koutis et al., 2011b]. We want to emphasize that the research on low stretch trees for general graphs is orthogonal to the main topic of this work. In our manuscript, we show in particular how to conduct efficient integration on arbitrary trees. Thus our work can be naturally combined with those algorithms to leverage all the above low stretch tree constructions for a better approximation of the graph’s metric.
The other class of method comes from the celebrated work of [Al-Mohy and Higham, 2011] and there are a number of extensions of this work [Kloster and Gleich, 2023, Al-Mohy and Higham, 2010, Moore, 2011, Moler and Van Loan, 2003, Auckenthaler et al., 2010].
Another class of methods is via sampling, where one samples a subset of a large matrix, which is then used to approximate the matrix-vector multiplication (i.e. Monte Carlo sampling) methods [Drineas et al., 2006, Drineas and Kannan, 2001, Acebron, 2019, Acebron et al., 2019, Benzi et al., 2017, Martinsson, 2019].
@@ Appendix C Topological Transformers
Algorithm 1 General Efficient Low-Rank Masked Attention from Choromanski et al. [2022]
@@
The central question in [Choromanski et al., 2022] was how to incorporate the masking in the linear attention as above. Note that in this case $\mathbf{A^{\prime}}$ is never materialized. Building on the work of [Luo et al., 2021], the authors [Choromanski et al., 2022] propose a general algorithm that efficiently implements masked linear attention.
@@ Appendix D Experimental Details and Additional Experiments
In this section, we provide additional details regarding the experimental setup and present additional results from our experiments. Our code is available at https://github.com/brcsomnath/FastTreeIntegrator. Specifically, we provide there the code for: (1) our algorithm leveraging IntegratorTree data structure (depicted in Fig 1), (2) adaptation to the Gromov-Wasserstein-type computation, (3) graph classification and (4) experiments on interpolation on meshes.
@@ D.1 Additional experiments for graph metric approximation with $f$-distance matrices
We also perform similar experiments for graph classification on the CUBES dataset Hanocka et al. [2019]. Specifically, we investigate how the polynomial degree affects the graph classification performance in Fig. 9 (left). We observe that increasing the polynomial degree improves the classification accuracy up to a certain degree. For the same dataset, we also compute the training loss for different polynomial degrees in Fig. 9 (right). Similarly, we observe that higher-degree rational functions achieve lower training loss for fitting the polynomial coefficients.
@@
Moreover, we benchmark FTFI on ModelNet10 [Wu et al., 2015], a dataset for 3D Point Cloud (PC) classification. For each PC, we create an $\epsilon$-neighborhood-graph and use FTFI for graph classification The Shortest Path kernel achieves an accuracy of $39.6\%$, whereas our FTFI with the degree-2 polynomial improves the accuracy to $44.2$% ($10$% relative improvement over the baseline), similar to the observation in 9.
@@ D.2 Integration of FTFI into GW-style algorithms
Wasserstein distance has found many uses in ML, particularly due to it’s principled approach to compare probability distributions. Gromov Wasserstein Mémoli [2011] discrepancy is an extension of Wasserstein distance to graph structured data, with a lot of downstream applications like graph clustering and classification. Inspired by the work of [Choromanski et al., 2023], we follow the exact same procedure in the integration of FTFI in the conditional gradient algorithm. The FTFI can be injected seamlessly in place of the Fast Matrix Multiplication (FMM) algorithms in Algorithm 2 and Algorithm 3 (see [Choromanski et al., 2023]).
Our method GW-FTFI run consistently $2$-$6$x faster than the baseline methods using the Shortest Path kernel, with no drop in accuracy in computing the associated costs (Figure 10). The plots shown are obtained by averaging over $10$ seeds and random trees of various sizes. For the baseline experiments, we use the implementation from the POT library [Flamary et al., 2021].
@@ D.3 Interpolation on Meshes
In the vertex normal prediction task in Section 4.2, we choose 40 meshes for 3D-printed objects with a wide range of size from the Thingi10K [Zhou and Jacobson, 2016] dataset with the File IDs:
@@ D.4 Additional Details on Graph Classification
We conduct graph classification experiments on a wide range of benchmark datasets. We report the dataset statistics for the graph classification datasets in Table 2. More details about the datasets are available in Morris et al. [2020]. To evaluate the performance of the different kernels, we employ the framework proposed by [Errica et al., 2020]. In particular, 10-fold cross-validation is used to obtain an estimate of the generalization performance of our method and the baseline method. We repeat this cross validation experiment 5 times to get a robust estimation and report the standard deviation for each setup.
To obtain graph features, we follow the approach presented in [de Lara and Pineau, 2018]. In this setting, we obtain the $k$-smallest eigenvalues from the approximated kernel from FTFI and forward these features to a random forest classifier for classification. For BGFI, we perform the same process obtaining the $k$-smallest eigenvalues from the exact shortest kernel. FTFI achieves similar performance to the BGFI while being significantly faster. We tune the hyperparameter $k$ independently for each method.
@@
Table 4: Comparison of FTFI with a broad range of graph kernel-based classification approaches. We observe that FTFI achieves performance similar to that of Exact SP, its exact counterpart, across almost all datasets. The baseline results have been compiled from Nikolentzos et al. [2021]. OOT and OOM indicate that the corresponding algorithm ran out of time or memory respectively.
@@ D.5 Additional details on experiments for Topological transformers
D.5.1 ImageNet
[…]
D.5.2 Places365
[…]
D.5.3 I-naturalist 2017
[…]
D.6 Video Vision Transformer
ViViT ([Arnab et al., 2021]) is a novel architecture that adapts the Vision Transformer (ViT) for video processing. It efficiently handles the spatiotemporal dimensions of video data by factorizing the input and applying attention mechanisms across both space and time. This allows ViViT to capture complex motion patterns and long-range dependencies in videos.
Applying FTFI with a topological masking mechanism to the ViViT architecture (factorized Transformer model variant, trained from scratch, as described in Arnab et al. [2021]) results in a $\mathbf{0.8}\%$ absolute improvement on the Kinetics dataset ([Kay et al., 2017]). The experimental setup follows Arnab et al. [2021]. To the best of our knowledge, this is the first application of Topological Transformers to video data.
@@ Appendix F Limitations
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We give detailed explanations of our contributions in the introduction (page 2).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations are clearly explained in Appendix F
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We introduce the notion of our algorithm Fast Tree Field Integrator in section 3. We describe the main algorithm in detail and introduce the technical (theoretical) results. The proofs of these results can be found in Appendix A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Training details to replicate each experiment are in the Appendix D.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide the code as well as details to run our experiments in Appendix D.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All details are presented in Sec 4 and Appendix D.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All experiments in the paper except for the ones using large Transformer models have been run multiple times using various random seeds and we report the relevant statistics. The experiments using Transformers are too expensive to run multiple times as the experiments are run on a huge dataset like ImageNet.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report the compute resources used in Appendix D.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: All authors have reviewed the NeurIPS code of ethics and the research conform to the code.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The broader impacts of our work is detailed in Appendix E.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper is theoretical in nature and we are not releasing any new models or data.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have properly cited all the papers that introduced various algorithms and data that are used in this work.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release the code for the main algorithm. The usage is detailed in the anonymous github repo.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not conduct any research that involves crowd sourcing or with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our paper does not involve crowd sourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.